# Corki: Enabling Real-time Embodied AI Robots via Algorithm-Architecture Co-Design

### v1

@@ Abstract.
Embodied AI robots have the potential to fundamentally improve the way human beings live and manufacture. Continued progress in the burgeoning field of using large language models to control robots depends critically on an efficient computing substrate. In particular, today’s computing systems for embodied AI robots are designed purely based on the interest of algorithm developers, where robot actions are divided into a discrete frame-basis. Such an execution pipeline creates high latency and energy consumption. This paper proposes Corki, an algorithm-architecture co-design framework for real-time embodied AI robot control. Our idea is to decouple LLM inference, robotic control and data communication in the embodied AI robots compute pipeline. Instead of predicting action for one single frame, Corki predicts the trajectory for the near future to reduce the frequency of LLM inference. The algorithm is coupled with a hardware that accelerates transforming trajectory into actual torque signals used to control robots and an execution pipeline that parallels data communication with computation. Corki largely reduces LLM inference frequency by up to 8.0×8.0$\times$8.0 ×, resulting in up to 3.6×3.6$\times$3.6 × speed up. The success rate improvement can be up to 17.3%. Code is provided for re-implementation. https://github.com/hyy0613/Corki
@@ 1. Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities in reasoning and long-term task planning (RT1, ; RT2, ; driess2023palm, ; huang2023voxposer, ; duan2022survey, ; song2023llm, ; liu2024robouniviewvisuallanguagemodelunified, ). Building upon the success of LLMs, the field of embodied AI, which employs LLMs to control robots interacting with the physical world, is increasingly recognized as a promising step towards achieving Artificial General Intelligence (AGI).
@@
The single most important difference between using LLMs for generating text and images versus integrating them as decision-making and planning modules within robotic pipelines lies in the hard real-time constraints imposed on robots (mei2006deployment, ; khatib1986real, ). Without real-time assurances, the applicability of embodied AI systems is severely limited to theoretical studies rather than real-world applications.
@@
However, current embodied AI systems struggle to meet real-time constraints. The fundamental reason lies in the execution model of embodied AI systems. To date, all embodied AI systems follow a sequential execution model that processes video input and generates robot actions in a frame-by-frame basis. Specifically, after warming up, the robots will start with a video sequence containing $N$ frames and a language instruction $i$. The LLM will predict the robot action tuple $(\Delta{x},\Delta{y},\Delta{z},...)$ based on the current input tuple $(Frame_{t-N},Frame_{t-N+1},...,Frame_{t},i)$, where $\Delta$ denotes the proposed robot movements and $Frame_{t}$ represents images within the video sequence. Upon executing the action, the robot captures a subsequent frame $Frame_{t+1}$ at the latest position. The next LLM inference then involves processing the updated input tuple $(Frame_{t-N+1},Frame_{t-N+2},...,Frame_{t+1},i)$.
@@
The current execution model is time-consuming due to two primary reasons. First, the sequential nature of each stage significantly contributes to the overall latency. Since most robots depend on high-end servers for LLM inference, the latency associated with the embodied AI systems is the cumulative effect of three distinct stages: LLM inference latency, robot action execution latency, and data communication latency. The sum of these latencies for each frame can add up to several hundred milliseconds. Second, all three stages have to happen for every frame, which further hurts the real-time constraints. We show this pipeline in Fig. 1a.
@@
(a) Current discrete execution pipeline of embodied AI systems, where every time a single next step action is predicted and all three stages happen for every frame.
@@
(b) Proposed continuous execution pipeline of embodied AI systems, where model predicts near future trajectory and pipelines communication latency with robot execution latency.
Fig. 1. Existing embodied AI systems pipeline and Corki pipeline.
@@
Today’s embodied AI pipeline is designed purely based on the convenience of algorithm designers as executing frame by frame sequentially is a traditional method in video processing algorithms, yet it violates a basic principle of robotic software design. That is, the front-end, responsible for perception and planning, does not inherently require real-time performance, whereas the back-end, which includes robot control algorithms, must operate in real-time. Critically, the unbalanced frequency requirements existing in robotic software stack allow us to decouple LLM inference, robotic control and data communication. After decoupling, we are able to reduce the front-end LLM inference rate, pipelining three stages and accelerating robotic control algorithms to achieve real-time performance in embodied AI applications.
@@
In this paper, we fundamentally change the execution pipeline of existing embodied AI systems to reduce the end-to-end latency. Firstly, at algorithm level, we depart from the conventional approach of predicting robot movement in the next frame discretely. Instead, we propose a novel embodied AI algorithm framework that is able to predict the trajectory of the robot for the near future. Unlike methods that focus on predicting only the immediate subsequent step, our algorithm accurately forecasts actions for multiple future steps. Thus, we significantly reduce the inference frequency of LLMs, saving both latency and energy.
Secondly, to accelerate the control process, we devise an accelerator capable of translating the trajectories predicted by LLMs into seamless and real-time control signals. The accelerator we design is tailored for task space computed torque control, with a customized data-flow accelerator, customized circuits and dedicated on-chip buffer design.
@@ Results.
We use a state-of-the-art embodied AI system, RoboFlamingo (li2023vision, ), as our baseline, and apply Corki on top of it. We show that Corki is able to achieve 3.6×3.6$\times$3.6 × speed up. The maximum success rate improvement is 17.3% higher than the baseline. The contribution of this paper is summarized as follows:
  • We observe that the existing embodied AI pipeline can not satisfy real-time constraints because of the sequential execution pipeline and balanced frequency of front-end video capture and back-end robot control. The high frequency of LLM inference and latency accumulation of every stage result in high latency.
  • We propose a new embodied AI algorithm framework to control robots by predicting future trajectories instead of the discrete movement of every frame.
  • We design an accelerator to smoothly transform the trajectory predicted by our models into robotic control signals in real time.
  • We design a new execution pipeline based on our proposed framework and hardware accelerator to hide communication latency between robot body and server.
@@
We organize our paper as follows. Sec. 2 introduces basic embodied AI system pipeline and motivates our paper. Sec. 3 introduces a new embodied AI algorithm framework that is able to predict continuous near-future trajectory of robots. Sec. 4 describes the proposed hardware accelerator for controlling robots given predicted trajectory and system pipeline design. Sec. 5 discusses the experimental methodology, followed by the evaluation results in Sec. 6. We discuss the related work in Sec. 7 and conclude our paper in Sec. 8.
@@ 2. Background and Motivation
We introduce the background of embodied AI systems (Sec. 2.1). We show that the execution pipeline of embodied AI systems is significantly different from previous utilization of LLMs and results in high end-to-end frame latency (Sec. 2.2).
@@ 2.1. Embodied AI System
Traditional robots typically depend on rule-based algorithms for decision-making and task planning, confining their utility to simple and predetermined scenarios. In contrast, the success of Large Language Models (LLMs) has spurred efforts to equip robots with advanced capabilities in reasoning and long-term planning. Such success boosts the emergence of applications that use LLMs for robot control, which has demonstrated notable advancements, particularly in enhancing the success rates of robots performing complex tasks in dynamic scenarios (li2023vision, ; wake2023gpt, ; hu2023toward, ; firoozi2023foundation, ).
@@
Embodied AI systems commonly employ a multi-modality Large Language Model  (lyu2023macaw, ; zhao2023bubogpt, ; mu2024embodiedgpt, ; huang2023chatgpt, ) as the planning module. This LLM seamlessly integrates language instruction inputs, such as ”put the blue mug on the table and bring me the red one,” with traditional sensor inputs in the robotic pipeline, including continuous videos, IMU signals, and point clouds (everett1995sensors, ; li2019common, ; santaera2015low, ). The LLM inference will generate the next actions for the robot body to perform based on current and recent observations along with the instructions.
Recently, embodied AI systems have demonstrated substantial potential to replace humans in various tasks. Google’s robotic transformer (RT1, ; RT2, ) has achieved an impressive success rate of over 75.0% on tasks including ”pick up objects”, ”open drawers”, and ”place objects into designated places” within domestic environments. RoboFlamingo (li2023vision, ), a recently proposed embodied AI framework, further elevates the success rate of a single task to over 89.5%.
@@ 2.2. Execution Pipeline and Performance Bottleneck
We use RoboFlamingo as an example of embodied AI systems to illustrate existing system pipelines. RoboFlamingo utilizes a vision-language model (VLM) to control a Franka Emika Panda robot arm with a parallel gripper (gaz2019dynamic, ), which in total has seven degrees of freedom. RoboFlamingo takes two forms of input, a language instruction and a video sequences containing 12 images. The model will predict the action of the robot arm’s end-effector within the next step. Equ. 1 describes the LLM inference process at frame $t$, where $F_{t}$ represents a single frame within a video sequence and $i$ denotes the language instruction. $\Delta x,\Delta y,\Delta z$ are the three-dimensional position change, $\Delta\alpha,\Delta\beta,\Delta\gamma$ are the three-dimensional rotation change and $g$ is the one-dimensional gripper status, which can be either open or closed.
After the model predicts the action, the robot arm will perform the action, moving itself to a new position. The control process on the robot translates movement information of the end-effector to the actual torque of each motor placed on the joints of the robot arm. A camera on the robot, usually placed on the gripper, will capture a new frame $F_{t+1}$ and send it back to the model to update the input frames. The next inference will happen on $(F_{t-10},F_{t-9},...,F_{t+1},i)$.
@@
Fig. 2. Latency and energy breakdown of RoboFlamingo framework.
Specifically, we analyze and characterize the execution pipeline of RoboFlamingo by breaking down the execution latency and show the results in Fig. 2. To get the results, we run LLM inference on a Nvidia V100 GPU, the robot control on an Intel 13th generation i7-13700 CPU and gather the data communication data using a Wi-Fi module.
Fig. 2a shows that even with a relatively small LLM (3 billion parameters) and a high-end GPU, the end-to-end frame latency of the embodied AI system can reach up to 249.4 ms, which directly contributes to a very low frame rate that does not satisfy real-time constraints. Among all three stages, LLM inference takes 76.9% of the execution time, robot control takes 4.1% and data communication takes 19.0%. Fig. 2b shows the energy breakdown. LLM inference still dominates with a 98.0% of the total energy, while robot control and data communication account for only 2.0%. Notice that the latency spent on control is low in the baseline system since the control frequency is set to match the front-end frame rate of 30 Hz. However, in real robotic systems, control usually has a much higher rate. Our characterization suggests that for each frame, to get a smooth trajectory, corresponding control latency can add up to 13.9% of the total latency.
@@
Detailed characterization data suggests the reasons contributing to slow execution of embodied AI robots are mainly twofold. First, the frame-by-frame sequential execution pipeline forces every action of the robot to undergo three stages: LLM inference, robot control and communication, and the latencies accumulate. Second, LLM inference, even with high-end GPU acceleration, is still extremely slow. The above reasons motivate this work.
From the perspective of a robotic system designer, the planning and decision making module does not need to match the high frequency of the control module. Trajectory is usually used as a bridge to eliminate the frequency mismatch. We use the same principle here.

### v5

# DaDu-Corki: Algorithm-Architecture Co-Design for Embodied AI-powered Robotic Manipulation

@@ Abstract.
Embodied AI robots have the potential to fundamentally improve the way human beings live and manufacture. Continued progress in the burgeoning field of using large language models to control robots depends critically on an efficient computing substrate, and this trend is strongly evident in manipulation tasks. In particular, today’s computing systems for embodied AI robots for manipulation tasks are designed purely based on the interest of algorithm developers, where robot actions are divided into a discrete frame basis. Such an execution pipeline creates high latency and energy consumption. This paper proposes Corki, an algorithm-architecture co-design framework for real-time embodied AI-powered robotic manipulation applications. We aim to decouple LLM inference, robotic control, and data communication in the embodied AI robots’ compute pipeline. Instead of predicting action for one single frame, Corki predicts the trajectory for the near future to reduce the frequency of LLM inference. The algorithm is coupled with a hardware that accelerates transforming trajectory into actual torque signals used to control robots and an execution pipeline that parallels data communication with computation. Corki largely reduces LLM inference frequency by up to 5.1×5.1$\times$5.1 ×, resulting in up to 5.9×5.9$\times$5.9 × speed up. The success rate improvement can be up to 13.9%.
Algorithm-architecture co-design, Embodied AI, Robotics, Hardware accelerator
journalyear: 2025copyright: ccconference: Proceedings of the 52nd Annual International Symposium on Computer Architecture; June 21–25, 2025; Tokyo, Japanbooktitle: Proceedings of the 52nd Annual International Symposium on Computer Architecture (ISCA ’25), June 21–25, 2025, Tokyo, Japandoi: 10.1145/3695053.3731099isbn: 979-8-4007-1261-6/2025/06ccs: Computing methodologies Robotic planningccs: Hardwareccs: Computer systems organization Real-time systems
@@ 1. Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities in reasoning and long-term task planning (Brohan et al., 2022, 2023; Driess et al., 2023; Huang et al., 2023a; Duan et al., 2022; Song et al., 2023; Liu et al., 2024). Building upon the success of LLMs, the field of embodied AI, which employs LLMs to control robots interacting with the physical world, is increasingly recognized as a promising step towards achieving Artificial General Intelligence (AGI).
@@
The single most important difference between using LLMs for generating text and images versus integrating them as decision-making and planning modules within robotic pipelines lies in the hard real-time constraints imposed on robots (Mei et al., 2006; Khatib, 1986). Without real-time assurances, the applicability of embodied AI systems is severely limited to theoretical studies rather than real-world applications.
In the robotics domain, there are two primary approaches to using LLMs. The first involves incorporating an LLM as the central decision-making module, where it decomposes complex tasks into fundamental robotic actions (Ahn et al., 2022; Liang et al., 2023; Vemprala et al., 2024). This high-level approach relies on the reasoning capabilities of LLMs and focuses on seamlessly integrating them into the robotics software pipeline. Here, the LLMs are typically very large in terms of parameters, can be single-modality, and are hosted on cloud servers (e.g., GPT-4). Given the low frequency of decision-making in robotic applications, the latency of communication and LLM inference can be tolerable. The second approach uses a smaller, usually multi-modality, LLM to guide basic robot control. This low-level case alters traditional program-based control and is highly sensitive to latency because the robot control module exercises a much higher frequency. Our work focuses on this second approach, and for the remainder of this paper, ’embodied AI systems’ will refer to this direction.
While using LLMs to control high degree-of-freedoms (DoF) robots such as humanoid robots may still be a distant goal, embodied AI algorithms have transformed manipulation tasks. Currently, the mainstream work flow for a robot to perform manipulation tasks without using embodied AI algorithms relies on efforts from expert programmers, where the programmer must specify the starting position, target position, and approach angles. Any change in the object’s location requires reprogramming. With advances in embodied AI, however, a robot arm can now manipulate different objects from any location on the table using only cameras and human instructions (Zhao et al., [n. d.]; Li et al., 2024; Octo Model Team et al., 2024; Wu et al., 2024; Collaboration et al., 2023; Cheang et al., 2024). This paper will focus on manipulation tasks.
@@
Current embodied AI systems struggle to meet real-time constraints. The fundamental reason lies in the execution model of embodied AI systems. To date, almost all embodied AI systems follow a sequential execution model that processes video input and generates robot actions on a frame-by-frame basis (Brohan et al., 2022, 2023; Octo Model Team et al., 2024; Kim et al., 2024; Collaboration et al., 2023). Specifically, after warming up, the robots will start with a video sequence containing $N$ frames and a language instruction $i$. The LLM will predict the robot action tuple $(\Delta{x},\Delta{y},\Delta{z},...)$ based on the current input tuple $(Frame_{t-N},Frame_{t-N+1},...,Frame_{t},i)$, where $\Delta$ denotes the proposed robot movements and $Frame_{t}$ represents images within the video sequence. Upon executing the action, the robot captures a subsequent frame $Frame_{t+1}$ at the latest position. The next LLM inference then involves processing the updated input tuple $(Frame_{t-N+1},Frame_{t-N+2},...,Frame_{t+1},i)$.
@@
The current execution model is time-consuming due to two primary reasons. First, the sequential nature of each stage significantly contributes to the overall latency. Since most robots depend on high-end servers for LLM inference, the latency associated with the embodied AI systems is the cumulative effect of three distinct stages: LLM inference latency, robot action execution latency, and data communication latency. The sum of these latencies for each frame can add up to several hundred milliseconds. Second, all three stages have to happen for every frame, further hurting the real-time constraints. We show this pipeline in Fig. 1a.
@@
(a) The current discrete execution pipeline of embodied AI systems, where every time a single next step action is predicted and all three stages happen for every frame.
@@
(b) Proposed continuous execution pipeline of embodied AI systems, where the model predicts near future trajectory and pipelines communication latency with robot execution latency.
Figure 1. Existing embodied AI systems pipeline and Corki pipeline.
@@
Today’s embodied AI pipeline is designed purely based on the convenience of algorithm designers, as executing frame by frame sequentially is a traditional method in video processing algorithms. Yet, it does not follow the design methodology in robotic domain and violates a basic principle of robotic software design that is widely adopted in robotic domain. The front-end, responsible for perception and planning, does not inherently require real-time performance. In contrast, the back-end, which includes robot control algorithms, must operate in real-time. The front-end and back-end can be aligned through a common intermediate representation: the robot’s movement trajectory. Critically, the unbalanced frequency requirements in the robotic software stack motivate us to decouple LLM inference, robotic control, and data communication. After decoupling, we can reduce the front-end LLM inference rate, pipelining three stages and accelerating robotic control algorithms to achieve real-time performance in embodied AI applications.
@@
In this paper, we fundamentally change the execution pipeline of existing embodied AI systems to reduce the end-to-end latency. Firstly, at the algorithm level, we depart from the conventional approach of predicting robot movement in the next frame discretely. Instead, we propose a novel embodied AI algorithm framework that is able to predict the trajectory of the robot for the near future. Unlike methods that focus on predicting only the immediate subsequent step, our algorithm accurately forecasts actions for multiple future steps. Thus, we significantly reduce the inference frequency of LLMs, saving both latency and energy.
Secondly, to accelerate the control process, we devise an accelerator capable of translating the trajectories predicted by LLMs into seamless and real-time control signals. The accelerator we design is tailored for task space computed torque control, with a customized data-flow accelerator, customized circuits, and dedicated on-chip buffer design. The most crucial architectural insight we find is that robotic control computations are performed at high frequencies, yet each control signal’s actual degree of change remains relatively low. This allows us to implement an application-specific approximate computing strategy. By analyzing the impact of individual joint movements on the control parameters, we can dynamically determine when to recompute these parameters and when to reuse previously computed values. This approximation significantly reduces computational overhead without compromising control accuracy.
@@ Results.
We use a state-of-the-art embodied AI system, RoboFlamingo (Li et al., 2024), as our baseline. Corki largely reduces LLM inference frequency by up to 5.1×5.1$\times$5.1 ×, resulting in up to 5.9×5.9$\times$5.9 × speed up. The success rate improvement can be up to 13.9%. The maximum success rate improvement is 17.3% higher than the baseline. The contribution of this paper is summarized as follows:
  • We observe that the existing embodied AI pipeline can not satisfy real-time constraints because currently the pipeline design is vision-centric, operating on a frame-by-frame basis, which results in high frame latencies.
  • We propose a new embodied AI algorithm framework from a robotic-centric angle to control robots by predicting future trajectories instead of the discrete movement of every frame, combined with a classic (non-LLM) high frequency controller.
  • We design a new execution pipeline based on our proposed framework to hide communication latency between the robot body and the server.
  • We design an accelerator to smoothly transform the trajectory predicted by our models into robotic control signals in real-time.
@@
We organize our paper as follows. Sec. 2 introduces basic embodied AI system pipeline and motivates our paper. Sec. 3 introduces a new embodied AI algorithm framework that is able to predict the continuous near-future trajectory of robots. Sec. 4 describes the proposed hardware accelerator for controlling robots given predicted trajectory and system pipeline design. Sec. 5 discusses the experimental methodology, followed by the evaluation results in Sec. 6. We discuss the related work in Sec. 7 and conclude our paper in Sec. 9.
@@ 2. Background and Motivation
We introduce the background of embodied AI systems (Sec. 2.1). We show that the execution pipeline of embodied AI systems is significantly different from the previous utilization of LLMs and results in high end-to-end frame latency (Sec. 2.2).
@@ 2.1. Embodied AI System
For manipulation tasks, traditional robots typically depend on rule-based algorithms for decision-making and task planning, confining their utility to simple and predetermined scenarios. In contrast, the success of Large Language Models (LLMs) has spurred efforts to equip robots with advanced reasoning and long-term planning capabilities. Such success boosts the emergence of applications that use LLMs for robot control, which has demonstrated notable advancements, particularly in enhancing the success rates of robots performing complex tasks in dynamic scenarios (Li et al., 2024; Wake et al., 2024; Hu et al., 2023; Firoozi et al., [n. d.]).
@@
Embodied AI systems commonly employ a multi-modality Large Language Model  (Lyu et al., 2023; Zhao et al., 2023; Mu et al., 2024; Huang et al., 2023b) as the planning module. This LLM seamlessly integrates language instruction inputs, such as ”put the blue mug on the table and bring me the red one,” with traditional sensor inputs in the robotic pipeline, including continuous videos, IMU signals, and point clouds (Everett, 1995; Li and Liu, 2019; Santaera et al., 2015). The LLM inference will generate the next actions for the robot body to perform based on current and recent observations along with the instructions.
Recently, embodied AI systems have demonstrated substantial potential to replace humans in various tasks. Google’s robotic transformer (Brohan et al., 2022, 2023) has achieved an impressive success rate of over 75.0% on tasks including ”pick up objects”, ”open drawers”, and ”place objects into designated places” within domestic environments. RoboFlamingo (Li et al., 2024), a recently proposed embodied AI framework, further elevates the success rate of a single task to over 89.5%.
@@ 2.2. Execution Pipeline and Performance Bottleneck
We use RoboFlamingo as an example of embodied AI systems to illustrate existing system pipelines. RoboFlamingo utilizes a vision-language model (VLM) to control a Franka Emika Panda robot arm with a parallel gripper (Gaz et al., 2019), which in total has seven degrees of freedom. RoboFlamingo takes two forms of input: a language instruction and a video sequence containing 12 images. The model will predict the action of the robot arm’s end-effector within the next step. Equ. 1 describes the LLM inference process at frame $t$, where $F_{t}$ represents a single frame within a video sequence and $i$ denotes the language instruction. $\Delta x,\Delta y,\Delta z$ are the three-dimensional position change, $\Delta\alpha,\Delta\beta,\Delta\gamma$ are the three-dimensional rotation change, and $g$ is the one-dimensional gripper status, which can be either open or closed.
After the model predicts the action, the robot arm will perform the action, moving itself to a new position. The control process on the robot translates the movement information of the end-effector to the actual torque of each motor placed on the joints of the robot arm. A camera on the robot, usually placed on the gripper, will capture a new frame $F_{t+1}$ and send it back to the model to update the input frames. The next inference will happen on $(F_{t-10},F_{t-9},...,F_{t+1},i)$.
@@
Figure 2. Latency and energy breakdown of RoboFlamingo.
[…]
Fig. 2a shows that even with a relatively small LLM (3 billion parameters) and a high-end GPU, the end-to-end frame latency of the embodied AI system can reach up to 249.4 ms, which directly contributes to a very low frame rate that does not satisfy real-time constraints. Among all three stages, LLM inference takes 72.7% of the execution time, robot control takes 9.9%, and data communication takes 17.4%. Fig. 2b shows the energy breakdown. LLM inference still dominates with 95.8% of the total energy, while robot control and data communication account for only 4.2%. Notice that the latency spent on control is low in the baseline system since the control frequency is set to match the front-end frame rate of 30 Hz. However, in real robotic systems, control usually has a much higher rate. Our characterization suggests that for each frame, to get a smooth trajectory, corresponding control latency can add up to 13.9% of the total latency.
@@
Detailed characterization data suggests that the reasons for the slow execution of embodied AI robots are mainly twofold. First, the frame-by-frame sequential execution pipeline forces every action of the robot to undergo three stages: LLM inference, robot control and communication, and the latencies accumulate. Second, LLM inference, even with high-end GPU acceleration, is still extremely slow.
The motivation for dedicated accelerators for control is clear. One of the key contributions of this work is reducing LLM inference frequency. However, our characterization shows that even if LLM inference latency were reduced to zero, the control frequency would still only reach 22.1 Hz, falling short of the real-time requirement (at least 30 Hz, 100 Hz preferable), while higher control rate allows the robot to better follow the predicted trajectory, increasing the safety of the control. Furthermore, control operations account for 39.7% of the total latency, with the rest latency spent on communication. The above reasons motivate the accelerator design. Note that we use the Intel i7-6770HQ CPU only because it is the onboard processor of our robot. We also tried to process the control algorithm with an Intel Core i7-13700 CPU and the corresponding frequency still can not meet real-time requirements.
From the perspective of a robotic system designer, the planning module does not need to match the high frequency of the control module. Trajectory is usually used as a bridge to eliminate the frequency mismatch. We apply the same principle.